# An Evolutionary Algorithm to Accelerate High-Dimensional Optimization

### v1

@@ 5.2. Implementation
QUASAR is implemented as a streamlined open-source Python package, designed for seamless integration into existing workflows.
The source code requires minimal dependencies, relying only on NumPy and SciPy. The open-source package, hdim-opt, is available at github.com/jgsoltes/hdim-opt for full use and reproducibility.
The package follows the syntax structures of SciPy’s optimization module, making it nearly one-to-one interchangeable with its Differential Evolution function.
@@ 7.1. Enhancements:
Crossover rate is currently defined by fitness rank, and may benefit from a probability-based approach similar to the other QUASAR components. Soft-max probabilities or similar methods to determine recombination probabilities may enhance genetic diversity.

### v2

@@ 5.2. Implementation
QUASAR is implemented as part of a streamlined open-source Python package, designed for seamless integration into existing workflows. The syntax follows the structure of SciPy’s optimization module, making it nearly one-to-one interchangeable with its Differential Evolution.
The source code requires minimal dependencies, relying only on NumPy and SciPy. The open-source package containing QUASAR, hdim-opt, is available via the Python Package Index (PyPI) using ’pip install hdim_opt’, or at github.com/jgsoltes/hdim-opt which includes the experimental notebooks for full reproducibility.
@@ 7.1. Enhancements:
Crossover rate is currently defined by fitness rank, and may benefit from a probability-based approach, similar to the other QUASAR components, via soft-max or similar methods.